# Arab Spring’s Impact on Science through the Lens of Scholarly Attention, Funding, and Migration

### v3

@@
However, there remains a notable absence of large-scale, systematic analyses that (1) identify who is studying which MENA countries, (2) trace how this attention varies across disciplines, and (3) reveal the evolution of this attention over time. Here, we address this gap by defining scholarly attention as the frequency with which countries are mentioned in article titles and abstracts, and take this as a first lens to understand the effect of the Arab Spring on the geographic distribution of research focus. Country naming practices have already proven effective in capturing global research patterns in domains ranging from public health  (?) and the social sciences  (?) to research on the United Nations’ Sustainable Development Goals (SDGs)  (?).
@@ Results
While the attention temporal trajectories differ across these countries, all of them received more annual scholarly attention in 2019 than in 2002 (Fig1c–e). Research on Egypt began accelerating after 2009, then rose above the global average between 2011 and 2015, dipped slightly in 2016, and then stabilized around the world’s mean (Fig. 1c). Attention towards Tunisia climbed steadily from 2005 until about 2010, after which it hovered at a stable plateau. In the civil war group (Fig. 1d), Yemen and Libya received minimal attention until 2009, followed by a surge that plateaued until 2012; Libya reached its peak in attention in 2013, while attention towards Yemen peaked in 2014, before both declined. Syria experienced a steady rise in attention, with a peak in 2016 probably corresponding to the rise of ISIS and associated humanitarian crises. Among the governmental changes group, attention towards Jordan and Morocco rose above the global median consistently; attention directed at Oman experienced a brief decline in 2012; attention towards Kuwait remained relatively flat; and Bahrain, despite a low overall volume, showed peaks in attention in 2008 and 2013, followed by growth after 2015.
@@
One possible explanation for the rise in scholarly attention is the increased research related to the Arab Spring. To assess how much of this attention is specifically driven by the Arab Spring’s aftermath, we investigate articles containing mentions of keywords related to the “Arab Spring” (e.g., Arab Uprisings, Arab Revolution, etc., see the Materials and Methods section for a comprehensive list and explanation of the methodology) in their title or abstract.
Figure 2: Scholarly attention due to Arab Spring-related research. (a), Country-specific scholarly attention related to research on the Arab Spring. Research focusing on the government overthrow (GO) group attracted the highest scholarly attention, followed by the civil war group (CW). (b), Proportion of Arab Spring-related research as part of total scholarly attention by target country. Syria and Libya, classified under the civil war group, exhibit the highest shares, whereas the governmental changes group shows minimal shares. The vertical dotted line represents the year 2011, marking the onset of the Arab Spring.
@@
The increase in scholarly attention of sources countries strongly correlates with their total scholarly output during the period (Pearson correlation: 0.684; Spearman correlation: 0.845, both with $p<0.0001$, see Fig. S4 in the Supplementary Materials) and with their average GDP during this period (Pearson correlation: 0.432; Spearman correlation 0.608, both with $p<0.0001$, see Fig. S5 in the Supplementary Materials). These two factors, GDP and scholarly output, possibly explain these disparities.
@@
To assess the impact of the Arab Spring on patterns of scholarly attention, we employ a difference-in-differences (DiD) statistical framework  (?). This approach, commonly used in econometrics and quantitative social sciences, involves employing observational data to simulate an experimental research design. It entails studying the differential effect of a treatment on a “treatment group” versus a “control group” in a natural experiment. This approach allows us to estimate the average “treatment ” effect of the Arab Spring by comparing changes over time in attention to our target countries, or in the three subgroups (government overthrow, civil war, and governmental changes) with those observed in either the rest of the world or a subset of MENA countries that remained relatively less affected during the same period (control groups). See the Materials and Methods section for a detailed explanation of the DiD approach and its application to our situation.
@@
Log Foreign Attention log GDP, log pop. Rest of world
@@
Table 1: Difference in differences estimates of the effect of Arab Spring on scholarly attention. Estimated coefficients from difference-in-differences regressions evaluating the post-treatment effects of the Arab Spring on log attention (total, domestic, foreign). Models control for log GDP, population, and research population, where applicable. Each colored column represents estimates from separate regressions, comparing the target countries with two control groups: the rest of the world and MENA countries less affected by the Arab Spring. Standard errors are in parentheses. ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. $\dagger$ denotes regressions where the parallel prior trend test fails to reject the null hypothesis of no difference in pre-Arab Spring trends at a 5% significance level.
Table 1 presents the difference-in-differences (DiD) analysis of scholarly attention. The columns labeled $\beta$ report the estimated effect of the 2011 Arab Spring on scholarly attention to the target countries. The analysis reveals that the Arab Spring had a significant positive effect ($p$ << 0.05) on scholarly attention towards affected countries compared to the rest of the world, and controlling for GDP, population, and researchers’ population. At the subgroup level, countries experiencing government overthrows and civil wars exhibit a positive treatment effect on total attention, with the civil war subgroup experiencing a stronger effect. In contrast, the effect for countries undergoing governmental changes is only statistically significant when compared to the rest of the MENA region, and it is negative. Further disaggregation by the source of attention offers additional insights. The effect on domestic attention is negative in the civil war subgroup compared to the rest of the world and the MENA region. In contrast, the effect is positive in the government overthrow subgroup, possibly reflecting the role of reform strategies. Regarding foreign attention, the effects are positive and statistically significant in the government overthrow and civil war subgroups, with the civil war subgroup exhibiting a notably higher effect. Conversely, the effect on foreign attention is negative in the governmental changes subgroup compared to the rest of the world and the MENA region.
@@
Collectively, these findings provide statistical evidence that the Arab Spring had an impact on scholarly attention toward the target countries and underscore the heterogeneity of this impact across the subgroups. While political instability, via government overthrows and civil wars, has sparked increased academic interest, governmental changes may dampen this engagement.
@@
To assess whether the Arab Spring significantly influenced funding patterns, we applied the same difference-in-differences framework used for scholarly attention (see Materials and Methods), using the log of total, domestic, and foreign funding as outcome variables while controlling for the log of population and GDP per capita. The results, presented in Table S5 of the Supplementary Materials, indicate that, for some regressions, the pre-treatment parallel trends assumption is not satisfied, and the model fit is generally poor. Nonetheless, the remaining regressions that still meet the required assumptions yield results that align with the patterns shown in Figure 7 (see Table S5 in the Supplementary Materials). The Civil War group shows a significant post-Arab Spring positive effect on funding relative to the rest of the world ( $p<0.01$). The effect is also positive on domestic funding for the governmental overthrow group and foreign funding for the civil war group (both significant at $p<0.01$ ). Together, these results underscore a post-uprising shift in both the level and source of funding, varying by political trajectory.
@@
Western countries, such as the UK, USA, France, and Germany, are also in the top 10 destinations, likely due to their strong economies and leadership in scientific research. The 10 destination countries are classified as HIC (high-income countries) and Upper-Middle-Income Countries (UMIC) in the Worldbank classification. There exists a weak, significant correlation between the change of migration to a country and its average GDP in the period (Spearman Correlation Coefficient: 0.33, $p<0.0001$, see Fig S11 in Supplementary Materials).
@@ Discussion
Our findings reveal that scholarly attention toward the ten countries affected by the Arab Spring has significantly increased in the post-Arab Spring period compared to the rest of the world, as measured by a difference-in-differences approach, and that this increase aligns with changes in funding and scholarly migration. Changes in scholarly attention cannot be attributed solely to the Arab Spring, as broader regional political, economic, and social dynamics likely contributed to shaping research trends. For instance, the long-term effects of the 2008 economic downturn, as well as the rise of ISIS in Syria and the Arab Winter, have also influenced science.
@@
We find varied patterns of foreign-to-domestic attention ratio across the target countries. For example, Egypt and Tunisia received more domestic attention than foreign attention before and after the Arab Spring. Countries experiencing civil wars received more foreign attention before the Arab Spring, and this imbalance grew even more post-Arab Spring, particularly in the social sciences and health sciences. Our findings show that shifts in scholarly attention are correlated with changes in research funding. Funding plays a crucial role in setting research agendas and influencing policy directions. While domestic funding has increased in countries experiencing government overthrows or governmental changes, foreign funding of research about the target countries has seen a more pronounced rise, particularly in nations affected by civil wars. Notably, the increase in funding has been most significant in the social sciences.
@@
These findings raise important questions about the optimal balance between domestic and foreign scholarly attention and its associated funding in fostering a sustainable research discourse. Domestic research and funding are crucial in strengthening national scientific productivity and research ecosystems, allowing local scholars to address context-specific challenges and build long-term research capacity. Meanwhile, foreign attention and funding enhance visibility and productivity and may facilitate integration into global academic networks. However, an overreliance on foreign attention may marginalize local voices in shaping discourse, potentially shifting knowledge production toward external and often temporary agendas rather than aligning with national priorities.
@@
Our findings also highlight the varied dynamics of scholarly migration: Egypt exhibited positive net migration rates, but Tunisia and Syria faced significant brain drain. The interplay between scholarly migration and scholarly attention may partly reflect institutional hiring practices, whereby researchers are employed due to their expertise in their country of origin  (?). In other cases, researchers may intentionally focus their work on their home country as an expression of national commitment, particularly in times of crisis, motivated by the hope of contributing to long-term development  (?). However, the mechanisms linking scholarly migration and patterns of scholarly attention remain underexplored, and further qualitative and quantitative research is needed to better understand these relationships.
@@
Our study has several limitations. First, it relies on Scopus-indexed publications, which primarily index English-language journals. As a result, contributions from local researchers publishing in non-English outlets may be underrepresented. Although accessing comprehensive local publication databases would offer valuable insight, doing so would require a multilingual country-detection algorithm and a scalable processing pipeline, both of which fall beyond the scope of this study. Furthermore, the quality, accessibility, and consistency of indexing in many local journals remain uneven, making it difficult to assess their reliability and integrate them systematically. Future research should address these gaps by developing tools to incorporate non-English and regionally published scholarship more effectively. Second, the difference-in-differences approach is limited by the difficulty of building control groups of similar countries unaffected by the Arab Spring. For example, countries in the remainder of the MENA region were influenced, to varying degrees, by the Arab Spring or its spillover effects from neighboring states. Nevertheless, in all analyses, we controlled for the size, GDP, and the size of the researcher population and ensured that the parallel trends assumption was satisfied; results that did not meet this criterion were explicitly reported. Finally, we measured scholarly attention by extracting mentions of countries in research articles. However, the absence of a country’s mention in a research paper does not necessarily indicate a lack of focus on that country, just as the presence of a country’s name does not automatically mean the study is centered on it. Despite these limitations, our study is the first large-scale investigation of the Arab Spring’s impact on scholarly attention and its relationship with funding and scholarly migration. Our framework enabled us to identify the significantly different trajectories taken by countries affected by the Arab Spring, as well as the foreign countries that emerged as key players in shaping academic research in the region.
This framework, comprising three lenses—scholarly attention, research funding, and scholarly migration—offers a scalable approach for constructing a comprehensive scientometric portrait of regional and global events. By combining it with empirical methods, such as quasi-experimental methods, it has wide applicability in the science of science  (?). For example, it can be applied to understand the impact of other conflicts, such as the Russia–Ukraine war, enabling a systematic investigation of how research priorities were reshaped: where scientific attention was directed, who funded the response, and how patterns of scholarly migration evolved. The framework can also be employed to assess the impact of political transitions, particularly changes in leadership or governance models, on both the international scholarly attention a country receives and the attention it directs outward. In this way, the framework provides a novel lens for observing how science policy agendas emerge, evolve, and circulate within the global research landscape.
@@
The temporal scope of our analysis spans from 2002 to 2019, covering the nine years preceding and following the onset of the Arab Spring in 2011, to capture the scholarly landscape surrounding this sociopolitical event. We chose not to analyze papers published after 2019 to avoid the disruptions caused by the COVID-19 pandemic, which significantly altered patterns of scientific mobility, collaboration, and scholarly attention. We focus on English articles to simplify the analysis of country mentions, avoiding the challenges of multilingual variations in country name spellings.
@@
Identifying scholarly attention to countries using country name extraction
@@
To extract mentions of countries or regions, we followed several steps. First, we removed copyright text from abstracts by deleting any content appearing after terms like ’@’ or ’Copyright (C)’, as such text may include the publisher’s country, which does not reflect the geographical focus of the research. Next, we used the Geotext library  (?) to identify and extract country names from titles and abstracts. Geotext employs country names, variations, and abbreviations from GeoNames, a comprehensive geographical database, to detect and classify mentions of cities and countries in textual data. Additionally, we used the pycountry library  (?) to map the extracted country names to their corresponding ISO Alpha-3 country codes, ensuring standardized representation. Then, we addressed specific cases to improve accuracy. Terms such as ‘US dollar,’ ‘New Mexico,’ ‘USD,’ ‘HK,’ and ‘Congo Red,’ and ‘Michael Jordan’, which contain geographical names in text but do not necessarily indicate the geographical focus of the research, were excluded from the list. These steps ensured that the extracted mentions accurately reflect scholarly attention to countries or regions.
@@
We also excluded demonyms (e.g., “Syrian,” “Egyptian”), as their presence does not necessarily indicate that the study focuses on the corresponding country. For example, research on the Syrian population in Lebanon or Germany may primarily address migration or diaspora dynamics, rather than Syria itself. Conversely, when research is conducted within a country, such as a study on “Egyptian goats” based in Egypt, the country is typically mentioned explicitly in the abstract or title, increasing the likelihood of accurate identification through direct country name matching.
@@
To enhance the algorithm’s accuracy, variations in country name spellings were incorporated into the dictionary and matched with regular expressions. For example, ‘U.S.,’ ‘United States,’ ‘United-States,’ ‘United States of America,’ ‘US’, and ‘U.S.A.’ were added to the dictionary for the USA.
We applied post-processing using specific criteria to enhance the accuracy of country identification, particularly in light of unique geopolitical circumstances. We excluded countries with populations under 1 million as of 2019, focusing our analysis on more populous nations.
@@
We encountered instances where country names could lead to false positives due to their similarity with terms commonly used in academic disciplines. For example, ”Jordan” shares its name with concepts in physics, mathematics, computer science, and engineering, named after Camille Jordan and Pascual Jordan (Including, but not limited to, Jordan Algebra, Jordan Normal Form, Jordan Measure, Jordan Matrix, and the Jordan–Wigner Transformation). We avoided including these subject areas to mitigate inaccuracies when studying the target countries. Additionally, we chose to exclude physics for another reason: the prevalence of large-scale (Mega) collaborations in this field, often involving hundreds of authors from multiple institutions and countries, makes it challenging to assess individual contributions to a single paper accurately  (?).
@@
Geopolitical sensitivities required additional adjustments. Cyprus and Macedonia were excluded because of their complex political divisions. Likewise, countries with ”Guinea” in their names—Guinea, Guinea-Bissau, Equatorial Guinea, and Papua New Guinea—were removed due to difficulties differentiating them using dictionary-based methods and their shared name with guinea pigs, a frequent subject in scientific and laboratory contexts. Similar exclusions were made for Ireland and Northern Ireland, the Republic of the Congo and the Democratic Republic of the Congo, as well as Sudan and South Sudan.
Countries that gained independence during our research period were also removed to avoid complications. Kosovo declared independence from Serbia in 2008, though its recognition remains disputed. Montenegro and Serbia became independent entities following the dissolution of the State Union of Serbia and Montenegro in 2006. East Timor (Timor-Leste) achieved full independence from Indonesia in 2002 after a UN-led transition.

### v4

@@
However, there remains a notable absence of large-scale, systematic analyses that (1) identify who is studying which MENA countries, (2) trace how this attention varies across disciplines, and (3) reveal the evolution of this attention over time. Here, we address this gap by defining scholarly attention as the frequency with which countries are mentioned in article titles and abstracts, and take this as a first lens to understand the effect of the Arab Spring on the geographic distribution of research focus. Country naming practices have already proven effective in capturing global research patterns in domains ranging from public health (?) and the social sciences (?, ?) to research on the United Nations’ Sustainable Development Goals (SDGs) (?).
@@ Results
While the attention temporal trajectories differ across these countries, all of them received more annual scholarly attention in 2019 than in 2002 (Fig. 1c–e). Research on Egypt began accelerating after 2009, then rose above the global average between 2011 and 2015, dipped slightly in 2016, and then stabilized around the world’s mean (Fig. 1c). Attention towards Tunisia climbed steadily from 2005 until about 2010, after which it hovered at a stable plateau. In the civil war group (Fig. 1d), Yemen and Libya received minimal attention until 2009, followed by a surge that plateaued until 2012; Libya reached its peak in attention in 2013, while attention towards Yemen peaked in 2014, before both declined. Syria experienced a steady rise in attention, with a peak in 2016 probably corresponding to the rise of ISIS and associated humanitarian crises. Among the governmental changes group, attention towards Jordan and Morocco rose above the global median consistently; attention directed at Oman experienced a brief decline in 2012; attention towards Kuwait remained relatively flat; and Bahrain, despite a low overall volume, showed peaks in attention in 2008 and 2013, followed by growth after 2015.
@@
One possible explanation for the rise in scholarly attention is the increased research related to the Arab Spring. To assess how much of this attention is driven explicitly by the Arab Spring’s aftermath, we investigate articles containing mentions of keywords related to the “Arab Spring” (e.g., Arab Uprisings, Arab Revolution, etc., see the Materials and Methods section for a comprehensive list and explanation of the methodology) in their title or abstract.
Arab Spring–related research is concentrated primarily within the Social Sciences category (refer to Supplementary Materials Fig. S8), in which the Social Sciences discipline accounts for 57.1% of the publications. The second-largest share belongs to Arts and Humanities (17.6%), followed by disciplines such as Economics, Econometrics and Finance (5.6%), and Business, Management and Accounting (3.7%), although their shares are considerably smaller.
Figure 2: Scholarly attention related to research on the Arab Spring. (a) Country-specific scholarly attention related to research on the Arab Spring. Research focusing on the government overthrow (GO) group attracted the highest scholarly attention, followed by the civil war group (CW). (b), Proportion of Arab Spring-related research as part of total scholarly attention by target country. Syria and Libya, classified under the civil war group, exhibit the highest shares, whereas the governmental changes group shows minimal shares. The vertical dotted line represents the year 2011, marking the onset of the Arab Spring.
@@
However, the temporal dynamics of scholarly attention in Arab Spring–related research not only vary across different political trajectories but also reveal distinct thematic emphases. To capture these differences, we applied BERTopic for topic modeling on the retrieved documents (see Supplementary Materials: Fig. S9, supplementary text, Table S3, and Table S4. Social media emerges as the most prominent theme across groups. In cases of government overthrow, the number of articles reached its peak in 2013. In the civil war, it reached its peak in 2015 and then decreased. For governmental changes, the rate has been growing since 2017. Social media played a dual role in the Arab Spring, facilitating collective action by enabling protest coordination while also shaping political debates and public discourse within a new media environment (?, ?, ?).
Topics such as Islamist parties, women’s rights, and financial and oil markets appeared predominantly in the government overthrow group, with only minimal presence in the other groups. By contrast, Turkey’s regional role and refugees were concentrated in the civil war group, reflecting how instability from civil conflicts reverberated into concerns over regional spillovers and broader geopolitical risks (?). Meanwhile, the theme of Gulf monarchies was characteristic of the governmental changes group, where references to Saudi Arabia and the Gulf Cooperation Council underscored the region’s geopolitical influence in shaping how the Arab Spring was studied within these countries.
@@
The increase in scholarly attention of sources countries strongly correlates with their total scholarly output during the period (Pearson correlation: 0.684; Spearman correlation: 0.845, both with $p<0.0001$, see Fig. S4 in the Supplementary Materials) and with their average GDP during this period (Pearson correlation: 0.432; Spearman correlation 0.608, both with $p<0.0001$, see Fig. S5 in the Supplementary Materials) and with pre–Arab Spring yearly average out-migration from target countries (Pearson correlation: 0.7813; Spearman correlation: 0.7453, both with $p<0.0001$, see Fig. S13 in the Supplementary Materials). These three factors–GDP, scholarly output, and pre–Arab Spring migration flows–possibly explain these disparities.
@@
To assess the impact of the Arab Spring on patterns of scholarly attention, we employ a difference-in-differences (DiD) statistical framework (?). This approach, commonly used in econometrics and quantitative social sciences, involves employing observational data to simulate an experimental research design. It entails studying the differential effect of a treatment on a “treatment group” versus a “control group” in a natural experiment. This approach allows us to estimate the average “treatment ” effect of the Arab Spring by comparing changes over time in attention to our target countries, or in the three subgroups (government overthrow, civil war, and governmental changes) with those observed in three control groups: the rest of the world, a subset of MENA countries that remained relatively less affected during the same period, and comparable countries selected through coarsened exact matching based on pre–Arab Spring averages of GDP per capita, population size, and researcher population (?). See the Materials and Methods section for a detailed explanation of the DiD approach and its application to our situation.
@@
CEM group 2
-0.1912
0.8038
0.0383
-0.3898∗∗∗,†
-0.2145
0.8015
Log Foreign Attention log GDP per capita, log pop. Rest of world
@@
Table 1: Difference in differences estimates of the effect of Arab Spring on scholarly attention. Estimated coefficients from difference-in-differences regressions evaluating the post-treatment effects of the Arab Spring on log attention (total, domestic, foreign). Models control for log GDP, population, and research population, where applicable. Each colored column represents estimates from separate regressions, comparing the target countries with three control groups: the rest of the world and MENA countries less affected by the Arab Spring and countries found through coarsened exact matching. Standard errors are in parentheses. ***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. $\dagger$ denotes regressions where the parallel prior trend test fails to reject the null hypothesis of no difference in pre-Arab Spring trends at a 5% significance level.
Table 1 presents the difference-in-differences (DiD) analysis of scholarly attention. The columns labeled $\beta$ report the estimated effect of the 2011 Arab Spring on scholarly attention to the target countries. The analysis reveals that the Arab Spring had a significant positive effect ($p$ << 0.05) on scholarly attention towards affected countries compared to the rest of the world, and controlling for GDP, population, and researchers’ population. Similar results ($p<0.1$) also hold when comparing to similar countries identified by coarsened exact matching (CEM).
At the subgroup level, countries experiencing government overthrows and civil wars exhibit a positive treatment effect on total attention, with the civil war subgroup experiencing a stronger effect. In contrast, the effect for countries undergoing governmental changes is only statistically significant when compared to the rest of the MENA region, and it is negative. When using a CEM-based control group, the estimates for government overthrow and civil war remain positive; however, the validity of these comparisons is limited by the failure to satisfy the parallel trends assumption ($p<0.05$) or test power. Similarly, the negative effect of governmental changes observed in the MENA comparison is not replicated in the CEM specification, further suggesting that results are sensitive to the choice of control group.
Further disaggregation by the source of attention offers additional insights. The effect on domestic attention is negative in the civil war subgroup compared to the rest of the world, the MENA region. In contrast, the effect is positive in the government overthrow subgroup, possibly reflecting the role of reform strategies. When using a CEM-based control group, the negative effect of civil wars remains, but the estimates again fail the parallel trends test, limiting their interpretability. For governmental changes, the CEM comparison replicates the negative effect observed in the broader control groups, this time significant at ($p<0.1$).
Regarding foreign attention, the effects are positive and statistically significant in the government overthrow and civil war subgroups, with the civil war subgroup exhibiting a notably higher effect. Conversely, the effect on foreign attention is negative in the governmental changes subgroup compared to the rest of the world and the MENA region. When the DiD analysis is conducted using the CEM control group, the government overthrow and civil war subgroups still show positive and statistically significant effects; however, because the estimates fail the parallel trends test, we refrain from drawing substantive conclusions from them.
@@
Collectively, these findings provide statistical evidence that the Arab Spring had an impact on scholarly attention toward the target countries, underscoring the heterogeneity of this impact across the subgroups. While political instability, via government overthrows and civil wars, has sparked increased academic interest, governmental changes may dampen this engagement.
@@
To assess whether the Arab Spring significantly influenced funding patterns, we applied the same difference-in-differences framework used for scholarly attention (see Materials and Methods), using the log of total, domestic, and foreign funding as outcome variables while controlling for the log of population and GDP per capita. The results, presented in Table S9 of the Supplementary Materials, indicate that, for some regressions, the pre-treatment parallel trends assumption is not satisfied, and the model fit is generally poor. Nonetheless, the remaining regressions that still meet the required assumptions yield results that align with the patterns shown in Figure 7 (see Table Table S9 in the Supplementary Materials). The Civil War group shows a significant post-Arab Spring positive effect on funding relative to the rest of the world ( $p<0.01$). The effect is also positive on domestic funding for the governmental overthrow group and foreign funding for the civil war group (both significant at $p<0.01$ ). Together, these results underscore a post-uprising shift in both the level and source of funding, varying by political trajectory.
@@
Western countries, such as the UK, USA, France, and Germany, are also in the top 10 destinations, likely due to their strong economies and leadership in scientific research. The 10 destination countries are classified as HIC (high-income countries) and Upper-Middle-Income Countries (UMIC) in the Worldbank classification. There exists a weak, significant correlation between the change of migration to a country and its average GDP in the period (Spearman Correlation Coefficient: 0.33, $p<0.0001$, see Fig. S12 in Supplementary Materials).
@@ Discussion
Our findings reveal that scholarly attention toward the ten countries affected by the Arab Spring has significantly increased in the post-Arab Spring period compared to the rest of the world, as measured by a difference-in-differences approach, and that this increase aligns with changes in funding and scholarly migration.
Changes in scholarly attention cannot be attributed solely to the Arab Spring, as broader global or regional political, economic, and social dynamics likely contributed to shaping research trends. For instance, the long-term effects of the 2008 economic downturn, as well as the Arab Winter (?) (a term used to describe the resurgence of authoritarianism, political violence, and civil conflict in the aftermath of the Arab Spring uprisings) and the rise of ISIS in Syria, have also influenced science. Topic analysis revealed themes such as social media, oil markets, women’s rights, refugees, and regional geopolitics, illustrating how different political trajectories shaped the evolution of research agendas of studying the Arab Spring itself.
@@
We find varied patterns of foreign-to-domestic attention ratio across the target countries. For example, Egypt and Tunisia received more domestic attention than foreign attention before and after the Arab Spring. Countries experiencing civil wars received more foreign attention before the Arab Spring, and this imbalance has grown even more pronounced post-Arab Spring, particularly in the social sciences and health sciences. Our findings show that shifts in scholarly attention are correlated with changes in research funding. Funding plays a crucial role in setting research agendas and influencing policy directions. While domestic funding has increased in countries experiencing government overthrows or governmental changes, foreign funding of research about the target countries has seen a more pronounced rise, particularly in nations affected by civil wars. Notably, the increase in funding has been most significant in the social sciences.
@@
These findings raise important questions about the optimal balance between domestic and foreign scholarly attention and its associated funding in fostering a sustainable research discourse. Domestic research and funding are crucial in strengthening national scientific productivity and research ecosystems, allowing local scholars to address context-specific challenges and build long-term research capacity. Meanwhile, foreign attention and funding enhance visibility and productivity and may facilitate integration into global academic networks. However, an overreliance on foreign attention may marginalize local voices in shaping discourse, potentially shifting knowledge production toward external and often temporary agendas rather than aligning with national priorities. Moreover, the temporary nature of such foreign engagements makes them susceptible to being overshadowed by events elsewhere or newer trends.
@@
Our findings also highlight the varied dynamics of scholarly migration: Egypt exhibited positive net migration rates, but Tunisia and Syria faced significant brain drain. The interplay between scholarly migration and scholarly attention may partly reflect institutional hiring practices, whereby researchers are employed due to their expertise in their country of origin (?). In other cases, researchers may intentionally focus their work on their home country as an expression of national commitment, particularly in times of crisis, motivated by the hope of contributing to long-term development (?). However, the mechanisms linking scholarly migration and patterns of scholarly attention remain underexplored, and further qualitative and quantitative research is needed to understand these relationships better.
@@
Our study has two main limitations. First, it relies on Scopus-indexed publications, which primarily index English-language journals. As a result, contributions from local researchers publishing in non-English outlets may be underrepresented. Although accessing comprehensive local publication databases would offer valuable insight, doing so would require a multilingual country-detection algorithm and a scalable processing pipeline, both of which fall beyond the scope of this study. Furthermore, the quality, accessibility, and consistency of indexing in many local journals remain uneven, making it difficult to assess their reliability and integrate them systematically. Future research should address these gaps by developing tools to incorporate non-English and regionally published scholarship more effectively. Second, we measured scholarly attention by extracting mentions of countries in research articles. However, the absence of a country’s mention in a research paper does not necessarily indicate a lack of focus on that country, just as the presence of a country’s name does not automatically mean the study is centered on it.
Despite these limitations, our study is the first large-scale investigation of the Arab Spring’s impact on scholarly attention and its relationship with funding and scholarly migration. Our framework enabled us to identify the significantly different trajectories taken by countries affected by the Arab Spring, as well as the foreign countries that emerged as key players in shaping academic research on the region.
This framework, comprising three lenses—scholarly attention, research funding, and scholarly migration—offers a scalable approach for constructing a comprehensive scientometric portrait of regional and global events. By combining it with empirical methods, such as quasi-experimental methods, it has wide applicability in the science of science (?).
For example, it can be applied to understand the impact of other conflicts, such as the Russia–Ukraine war, enabling a systematic investigation of how research priorities were reshaped: where scientific attention was directed, who funded the response, and how patterns of scholarly migration evolved. The framework can also be employed to assess the impact of political transitions, particularly changes in leadership or governance models, on both the international scholarly attention a country receives and the attention it directs outward. In this way, the framework provides a novel lens for observing how science policy agendas emerge, evolve, and circulate within the global research landscape.
@@
The temporal scope of our analysis spans from 2002 to 2019, covering the nine years preceding and following the onset of the Arab Spring in 2011, to capture the scholarly landscape surrounding this sociopolitical event. We chose not to analyze papers published after 2019 to avoid the disruptions caused by the COVID-19 pandemic, which significantly altered patterns of scientific mobility, collaboration, and scholarly attention. We focus on articles with English titles and abstracts to simplify the analysis of country mentions, avoiding the challenges of multilingual variations in country name spellings.
@@
Identifying scholarly attention to countries with country name extraction
@@
To extract mentions of countries or regions, we followed several steps. First, we removed copyright text from abstracts by deleting any content appearing after terms like ’@’ or ’Copyright (C)’, as such text may include the publisher’s country, which does not reflect the geographical focus of the research. Next, we used the Geotext library (?) to identify and extract country names from titles and abstracts. Geotext employs country names, variations, and abbreviations from GeoNames, a comprehensive geographical database, to detect and classify mentions of cities and countries in textual data. Additionally, we used the pycountry library (?) to map the extracted country names to their corresponding ISO Alpha-3 country codes, ensuring standardized representation. Then, we addressed specific cases to improve accuracy. Terms such as ‘US dollar,’ ‘New Mexico,’ ‘USD,’ ‘HK,’ and ‘Congo Red,’ and ‘Michael Jordan, ’ which contain geographical names in text but do not necessarily indicate the geographical focus of the research, were excluded from the list. These steps ensured that the extracted mentions accurately reflect scholarly attention to countries or regions.
@@
We also excluded demonyms (e.g., “Syrian,” “Egyptian”), as their presence does not necessarily indicate that the study focuses on the corresponding country. For example, research on the Syrian population in Lebanon or Germany may primarily address migration or diaspora dynamics, rather than Syria itself. In addition, many demonyms overlap with language names (e.g., French, Vietnamese, Georgian), creating further ambiguity, as such terms may refer to the language rather than the people or the country. Conversely, when research is conducted within a country, such as a study on “Egyptian goats” based in Egypt, the country is typically mentioned explicitly in the abstract or title, increasing the likelihood of accurate identification with only direct country name matching.
@@
To enhance the algorithm’s accuracy, variations in country name spellings were incorporated into the dictionary and matched with regular expressions. For example, for the USA, ‘U.S.,’ ‘United States,’ ‘United-States,’ ‘United States of America,’ ‘US’, and ‘U.S.A.’ were added to the dictionary.
@@
We applied post-processing using specific criteria to enhance the accuracy of country identification. First, we excluded countries with populations under 1 million as of 2019, focusing our analysis on more populous nations. Second, we encountered instances where country names could lead to false positives due to their similarity with terms commonly used in academic disciplines. For example, ”Jordan” shares its name with concepts in physics, mathematics, computer science, and engineering, named after Camille Jordan and Pascual Jordan (Including, but not limited to, Jordan Algebra, Jordan Normal Form, Jordan Measure, Jordan Matrix, and the Jordan–Wigner Transformation). We avoided including these subject areas to mitigate inaccuracies when studying the target countries. Third, we chose to exclude physics for another reason: the prevalence of large-scale (Mega) collaborations in this field, often involving hundreds of authors from multiple institutions and countries, makes it challenging to assess individual contributions to a single paper accurately (?).
@@
Additionally, geopolitical sensitivities required additional adjustments. Cyprus and Macedonia were excluded because of their complex political divisions. Likewise, countries with ”Guinea” in their names, Guinea, Guinea-Bissau, Equatorial Guinea, and Papua New Guinea, were removed due to difficulties differentiating them using dictionary-based methods and their shared name with guinea pigs, a frequent subject in scientific and laboratory contexts. Similar exclusions were made for Ireland and Northern Ireland, the Republic of the Congo and the Democratic Republic of the Congo, as well as Sudan and South Sudan. Countries that gained independence during our research period were also removed to avoid complications. Kosovo declared independence from Serbia in 2008, though its recognition remains disputed. Montenegro and Serbia became independent entities following the dissolution of the State Union of Serbia and Montenegro in 2006. East Timor (Timor-Leste) achieved full independence from Indonesia in 2002 after a UN-led transition.